The stream function for a two dimensional flow is given by ψ = 8xy, calculate the velocity at the point p(4, 5). Find the velocity potential function φ.
1
Expert's answer
2020-12-10T20:25:24-0500

In velocity we can write 


"u=\dfrac{\partial\psi}{\partial y}, v=-\dfrac{\partial\psi}{\partial x}"

Given "\psi=8xy"

"u=\dfrac{\partial\psi}{\partial y}=8x, v=-\dfrac{\partial\psi}{\partial x}=-8y"

Calculate the velocity at the point p(4, 5)


"u=32, v=-40"

"|velocity|=\sqrt{u^2+v^2 }=\sqrt{(32)^2+(-40)^2}=8\sqrt{41}"


"u=\dfrac{\partial\varphi}{\partial x}, v=\frac{\partial\varphi}{\partial y}"

"u=\dfrac{\partial\varphi}{\partial x}=8x"

Integrate with respect to "x"


"\varphi=4x^2+g(y)"

"\frac{\partial\varphi}{\partial y}=\frac{dg}{d y}=-8y"

Integrate with respect to "y"


"g(y)=-4y^2+C"

Then


"\varphi(x,y)=4x^2-4y^2+C"